### v2

# Denis Roio Dyne.org &Rebecca Selvaggini

Abstract
It is of critical importance to design digital identity systems that ensure the privacy of citizens as well as protecting them from issuer corruption. We aim to solve this issue and propose a method for selective disclosure and privacy preserving revocation of digital credentials, using the unique homomorphic characteristics of second order Elliptic Curves and Boneh-Lynn-Shacham (BLS) signatures. Our approach ensures that users can selectively reveal credentials signed by a certain issuer, which can be interactively revoked by a quorum of other agreeing issuers without revealing the identity of users. Our goal is to protect users from issuer corruption by requiring collective agreement among multiple revocation issuers.
1 Introduction
Digital identity systems implement credential issuance and presentation mechanisms so that a person (holder) can voluntarily disclose his or her own acquired skills, professed attributes, or completed accomplishments. Credentials are signed by issuer authorities and encapsulated within various forms of digital proofs to be held in digital wallets, empowering individuals to reveal only chosen details to designated recipients, to limit data exposure and permit a user-controlled release of information.
Such systems are known as selective disclosure and they enhance users privacy by allowing data minimization when proving credentials.
1.1 State of the art
Selective disclosures are being used by nation states across the world in their next generation identity systems, for instance EIDAS2.0 in Europe where the European Digital Identity Wallet Architecture and Reference Framework[1] mandates the use of SD-JWT[2] and mDOC [3]. SD-JWT adopts for its cryptography Hash-Based Message Authentication Codes (HMAC) to generate proofs: such presentations can be traced and recent critical feedback on the EUDI ARF[4] details this problem.
In North America the efforts concentrate on the adoption of the BBS+ algorithm[5] leveraging its Zero Knowledge Proof properties and applied to W3C Verifiable Credentials[6] to obtain an higher degree of privacy by making every disclosure unlinkable.
1.2 Threats considered
The different choices in data formats in these two approaches is irrelevant in relation with cryptography, any choice between JSON Web Tokens or W3C Verifiable Credentials does not impact the privacy level. But the cryptography adopted determines the adequacy of a solution to face three important threats that can render an algorithm unsuitable to be used in real world situations.
Linkability
The EUDI-ARF standard dictates that credentials issued to a holder, can be presented (in the form of a verifiable presentation) to a relying party in order to have one or more attributes verified. Every verifiable presentation includes one or more HMAC(s), formatted in SD-JWT: the HMACs are identical each time a verifiable presentation is produced from a certain credential. This makes possible for colluding relying parties, or to malicious actors, to trace a holder’s identity by collecting, exchanging and confronting verifiable presentations (linkability). This threat appears to be well mitigated by BBS+ through its Zero Knowledge Proof implementation.
Privacy breach of revocation lists
We believe that anonymous revocations are a condicio sine qua non to guarantee a sufficient level of privacy in digital identities and credentials. There is no privacy-preserving revocation system designed, either in EUDI-ARF, or in W3C-VC and BBS+. In case the choice of strategy for revocation is left open to developers, the risk for major privacy breaches may occur, for example with the adoption of public status lists [7]. The hypotetical use of a certificate status lists (CRL) presents issues related primarily to privacy [8], because sensitive information about holders leaks from the list. This problem is partially mitigated by expiration dates, in cases where credentials can be short-lived (typically less important credentials), but not applicable with digital identification documents such as ID, driving license, passport and social security numbers, which typically have longer or no expiration time. Future plans for the national standards we are observing include the adoption of “Bitstring” status lists [9] which may grant a degree of privacy. Our approach is that of designing a privacy-preserving revocation mechanism to remove the leak of holder’s information and allow to delegate the governance of revocations to multiple revocation issuers which may be different from the credential issuer.
Revocation issuer corruption
If the choice of interactive revocation is left to a single issuer, one may unilaterally choose to revoke credentials, without being subject to revision or having to seek consensus with a quorum of issuers. This situation leads to security issues in case Issuers are corrupted and make a weaponized use of digital revocations to persecute engaged individuals. Such a condition becomes a real concern for journalists or activists living under dictatorial regimes that may arbitrarily revoke their credentials, or even ID cards and passports. Similarly, a security breach of an issuer service, would result in similar threats. We mitigate this risk by introducing the possibility for threshold issuance of revocation keys and by separating the responsibility of revocation issuance and credential issuance.
Feature Comparison
UP UR TR
SD-BLS no yes yes
SD-JWT no wip no
BBS+ yes wip no
Table 1: feature comparison
We briefly round up on feature differences between the named selective disclosure cryptographic schemes, as shown in table 1, mainly distinguishing between three fundamental features:
  • UP: Unlinkable Presentation
  • UR: Unlinkable Revocation
  • TR: Threshold Revocation
Where wip is mentioned meaning work in progress on adoption of bitstring status lists for unlinkable revocations.
2 Overview
2.1 Key contributions
The cryptographic scheme described in this paper, named SD-BLS for brevity, implements some of the properties of the SD-JWT scheme (which can be easily encoded in mDOC format) and proposes a novel cryptographic approach to similar data structures. Furthermore SD-BLS proposes novel anonymous cryptographic revocation flow for verifiable credentials, that aims to solve the privacy issues posed by status and revocation lists.
Selective disclosure
Similarly to the SD-JWT and mDOC formats, SD-BLS produces an array of claims: the elements of the array are individually signed by the issuer. In SD-BLS the signature(s) replace the HMAC. Using the homomorphic properties of this curve SD-BLS enables the holder to selectively disclose only certain issuer signed claims and produce a presentation that minimizes private information given to verifiers.
Anonymous cryptographic revocation
SD-BLS proposes a novel approach to credential revocation: the data published by the revocation issuer will produce cryptographic material that contains no information about the credential holders. The cryptographic revocation material allows anyone to verify if an SD-BLS proof produced by a credential holder has been revoked by one or multiple issuers. The unlinkability, and thus anonymity of the cryptographic revocation, allows the revocation issuer to share revocations in public and allows anyone to verify if credentials have been revoked.
2.2 Applications
In this section we present some applications and use cases for digital identity and credentials, that could benefit from using the SD-BLS scheme.
Digital identity
The focus of the EUDI-ARF specifications is identity documents: it defines mechanisms and data structures to issue a Personal Identification (PID) as well as digital driving licenses. Similarly, the US government is experimenting with W3C-VC and mDOC for cross-states interoperable driving licenses. SD-BLS data format is similar to SD-JWT and mDOC, offering selective disclosure and anonymous revocation.
Academic credentials
Diploma and academic credentials are among the core offerings of EBSI as well as a primary research target of the W3C VC working group.
KYC/AML
We are unaware of standardization efforts for interoperable credentials in the fields of "Know Your Customer" (KYC) and Anti Money-Laundering (AML) certifications. We are aware of solution providers experimenting with W3C-VC for AML applications and believe SD-BLS can greatly improve the governance of credential revocation, which is a critical component for this use case.
Generic light credentials
As the digital identity and verifiable credential technologies are maturing, they are being considered for usage in less privacy concerning applications, such subscriptions and membership and fidelity cards.
Verifiable credentials on Blockchain
SD-BLS can be used with blockchain-based smart-contracts to activate certain functions:
  • A verifiable presentation can be used by a smart-contract to verify if one or more holder’s claim match the requirement needed to process a transaction.
  • Issuers can publish their cryptographic revocation lists on chain, allowing smart-contracts to verify the status of a credential.
  • The issuer’s public keys can also be published on-chain, although this does not represent a novelty.
3 Implementation
In this section we will provide a detailed description of the algorithm we propose for selective disclosure and unlinkable revocation using BLS signatures.
3.1 Notations and assumptions
We will adopt the following notations:
  • $\mathbb{F}_{p}$ is the prime finite field with $p$ elements (i.e. of prime order $p$);
  • $E$ denotes the (additive) group of points of the curve BLS12-381 [10] which can be described with the Weierstrass form $y^{2}=x^{3}+16$;
  • $E_{T}$ represents instead the group of points of the twisted curve of BLS12-381, with embedding degree $k=12$. The order of this group is the same of that of $E$;
We also require defining the notion of a cryptographic pairing. That is defined as a function $e:\mathbb{G}_{2}\times\mathbb{G}_{1}\to\mathbb{G}_{T}$, where $\mathbb{G}_{1},\mathbb{G}_{2}$ and $\mathbb{G}_{T}$ are all groups of same order $n$, such that satisfies the following properties:
  • i.
    Bilinearity, i.e. given $P_{1},Q_{1}\in\mathbb{G}_{1}$ and $P_{2},Q_{2}\in\mathbb{G}_{2}$, we have
    $e(P_{2},P_{1}+Q_{1})=e(P_{2},P_{1})\cdot e(P_{2},Q_{1})$
    $e(P_{2}+Q_{2},P_{1})=e(P_{2},P_{1})\cdot e(Q_{2},P_{1})$
  • ii.
    Non-degeneracy, meaning that for all $g_{1}\in\mathbb{G}_{1},g_{2}\in\mathbb{G}_{2}$, $e(g_{2},g_{1})\neq 1_{\mathbb{G}_{T}}$, the identity element of the group $\mathbb{G}_{T}$;
  • iii.
    Efficiency, so that the map $e$ is easy to compute;
  • iv.
    $\mathbb{G}_{1}\neq\mathbb{G}_{2}$, and moreover, that there exist no efficient homomorphism between $\mathbb{G}_{1}$ and $\mathbb{G}_{2}$.
For the purpose of our protocol we will have $\mathbb{G}_{1}=E$ and $\mathbb{G}_{2}=E_{T}$, and $\mathbb{G}_{T}\subset\mathbb{F}_{p^{12}}$ is the subgroup containing the $n$-th roots of unity, where $n$ is the order of the groups $E$ and $E_{T}$. Instead $e:E_{T}\times E\to\mathbb{G}_{T}$ is the Miller pairing, which in our work is encoded as the method miller(ECP2 P, ECP Q).
3.2 Issuance
As for other well known algorithms BLS signing will work following three main steps:
  • Key Generation phase. For an issuer who wants to sign a credential $m$, a secret key $sk$ is a random number chosen uniformly in $\mathbb{F}_{n}$, where $n$ is the order of the groups $\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T}$. The corresponding public key $pk$ is the element $sk\cdot G_{2}\in E_{T}$;
  • Signing phase. The credential $m$ is first hashed into the point $U\in E$, which in our scheme is done by the method hashtopoint; the related signature is then given by $\sigma=sk\cdot U$;
  • Verification phase. For an other user that wants to verify the authenticity and the integrity of the message $m$, it needs to
    • 1.
      parse $m,pk$ and $\sigma$
    • 2.
      hash the message $m$ into the point $U$ and then check if the following identity holds,
      $e(pk,U)=e(G_{2},\sigma)$
    If verification passes it means that $\sigma$ is a valid signature for $m$.
  • Proof of correctness By using the definitions of the elements involved and exploiting the property of the pairing $e$ we have
    $\begin{split}e(pk,U)&=e(sk\cdot G_{2},U)\\ &=e(G_{2},U)^{sk}\\ &=e(G_{2},sk\cdot U)\\ &=e(G_{2},\sigma)\end{split}$
BLS signatures also support aggregation: it is possible to aggregate a collection of multiple signatures $\sigma_{i}$ (each one related to a different message $m_{i}$) into a singular new object $\sigma$, that can be validated using the respective public keys $pk_{i}$ in a suitable way.
Since $\sigma_{i}\in G_{1}\forall i$, the algorithm has an homomorphic property. We exploit this property to add a revocation signature into the signed credential.
We create a new secret key $rev$ with public key $r$ and signature of the claim $\sigma_{rev}$ to add to the issuer signature:
$\begin{split}r&=rev\cdot G_{2}\\ \sigma_{rev}&=rev\cdot U\\ \sigma&=sk\cdot U+\sigma_{rev}\\ \end{split}$
The set of all the signed claims will be:
$\mathcal{C}=\big{\{}\{id,r,\sigma\}:id\in claims\big{\}}$
At the end of this phase the holder is sent the signed claims to be stored in a private wallet, while the issuer stores a list of revocations, associated with an identifier of the holder and the specific credential, into a private database that can be used later to issue revocations.
3.3 Presentation
When the holder needs to disclose a particular claim $id$ it is sufficient to send the signed claim set $\{id,r,\sigma\}$.
3.4 Verification
Credential verification is made by checking the presented issuer’s signature. In order to do that the credential issuer’s public key must be added to the revocation public key.
As explained in section 3.3 we can consider the credential proof as a collection of tables of the following form:
$c_{id}=\{id,r,\sigma\}$
where $\sigma$ and $r$ are respectively the signature of the string $id$ and the revocation public key. We can check the validity of the presented claim computing the key
$pk=A.pk+r$
and verify the bls signature $\sigma$.
3.4.1 Proof of correctness
The signature $\sigma$ is given by
$\begin{split}\sigma&=sign(A.sk,id)+sign(rev,id)\\ &=A.skU+revU\end{split}$
where $U$ is the mapping of the string $id$ in the group $G_{1}$.
Recalling the verification formula, it holds that:
$\begin{split}e(pk,U)&=e(A.skG_{2}+revG_{2},U)\\ &=e(A.skG_{2},U)\cdot e(revG_{2},U)\\ &=e(G_{2},A.skU)\cdot e(G_{2},revU)\\ &=e(G_{2},A.skU+revU)\\ &=e(G_{2},\sigma)\\ \end{split}$
where each equality holds for the bilinearity of the Miller loop.
This concludes the first verification phase. If the given presentation is valid, then the verifier should proceed to check the revocation status of the credential as explained below.
3.5 Revocation
To control if a revocation has been emitted for any credential being verified, we update the revocation list from the issuer. A revocation list can be publicly distributed since revocation keys do not provide any information on the identity of holders.
Given an element $c_{id}=\{id,r,\sigma\}$ of the credential presentation and the Issuer public key $A.pk$, if the claim id is present in the revocation list, we can take the corresponding revocation private key $rev$.
We can verify the validity of the revocation by checking if the revocation public key presented by the holder match:
$r=G_{2}\cdot rev$
This step should follow the signature verification and is sufficient to guarantee the revocation status.
If a dishonest holder provides a wrong $r$, then the verification of the signature will fail, and the given credential should be considered invalid.
If it happens that the signature verification is successful and that the revocation key does not match, one can conclude that the credential is not revoked.
3.6 Threshold Revocation
In order to split responsibilities over interactive revocation we introduce a threshold over the revocation key, plus we split the functionality of credential issuance from that of revocation issuance, now operated by different peers.
This is implemented via an interactive process facilitated by a third party trustee, a revocation dealer, who should be:
  • never entitled to publish revocations
  • connected to credential issuers to complete any credential signature
  • never informed about the identity of credential holders
  • regularly connected to revocation issuers to distribute shares
Such a revocation dealer will be in contact with credential issuers for the signature of credentials: it will create the $rev$ revocation key while concealing it from them. The trustee then proceeds creating the $\sigma_{rev}=rev\cdot U$ revocation signature and the $r=rev\cdot G_{2}$ public key as detailed in the issuance phase (3.2). Finally the revocation trustee will distribute $\sigma_{rev}$ and $r$ to issuers for further processing into the final signed credential.
The revocation trustee (dealer) will then proceed asynchronously to split the revocation key into shares using a public verifiable secret sharing (PVSS [11]) implementation and distribute these shares to all revocation issuers. In order to issue a revocation, a configurable quorum of peers among the revocation issuers will need to reconstruct the secret revocation key and publish it.
This process separates responsibilities between the credential issuer and the revocation issuers, delegating to the revocation issuers the possibility to revoke a credential interactively through a collective process.
The collection of shares can be done asynchronously and is provable. The dealer should publish proofs of knowledge of each revocation share, proving their creation and authenticity. Revocation issuers can also use the dealer proofs to verify the validity of the shares received without revealing them. Such revocation proofs will also be useful when revocation issuers will reconstruct a revocation, since they can refer to them to prove their shares are authentic without revealing their content.
4 Benchmarks
We implemented the flows for credential issuance, presentation, verification and revocation for lab tests using Zenroom  [^1] Zenroom home: https://zenroom.org, a secure isolated execution environment implementing advanced cryptography transformations. The reference implementation for this paper is published on a public repository  [^2] SD-BLS github repository: https://github.com/dyne/sd-bls. All benchmarks were executed on a 6th gen. Intel PC running tests on a single i7 3.40GHz core and making no use of hardware acceleration.
Lab measurements on a growing number of claims show that issuance is less computation heavy than verification, as shown in figure 1.
Figure 1: Speed comparison of issue and verify
Based on our benchmarks, the resulting data objects sizes, excluding the signed material which can be either of variable length or fixed length hashes, are:
  • Proof: 145 Bytes
  • Revocation: 32 Bytes
The computational cost grows for the verification phase with the presence of a cryptographic revocations list. We assume that revocations are grouped by claim type or credential structure, for instance all revocations of a certain common claim (e.g.: "is above 18" or "is a Danish citizen") for any holder are published in the same list: this will not degrade the level of privacy and will group some revocations in smaller lists.
Lab measurements of the time taken by a single proof verification process to operate on a growing number of revocations demonstrate that there is a linear growth, as shown in figure 2.
Figure 2: Speed of verification of a claim over multiple revocations
The threshold operated for revocation issuers consists of a verifiable secret sharing implementation supporting a configurable total and quorum of peers. Our implementation shows very good performance on reconstruction, which is the most speed-sensitive operation in scenarios where responsive revocation process is required. The process of reconstruction can be easily scaled for asynchronous consensus on a micro-service swarm architecture and verified on blockchain.
Figure 3: Speed of creation and reconstruction of shares among multiple peers
5 Conclusion
5.1 Security considerations
The revocations database is privacy and corruption sensitive (by our previous definition of issuer corruption) and it should be securely stored by each Issuer. This is mitigated by the adoption of threshold revocation. When using threshold revocation there is still the need for a revocation trustee who may be a single point of failure if allowed to publish revocation keys dishonestly collected during the process of credential issuance.
BLS signatures and the proof system obtained with credentials are considered secure by assuming the existence of random oracles [12], together with the decisional Diffie-Hellman Problem (DDH) [13], the external Diffie-Hellman Problem (XDH), and with the Lysyanskaya-Rivest-Sahai-Wol Problem (LRSW) [14], which are connected to the Discrete Logarithm.
The SD-BLS implementation we are presenting in this paper is demonstrated using the BLS12-381 curve [10] also adopted by ETH2.0. Debating the choice of BLS12-381 is beyond the scope of this paper, but is worth mentioning that we can easily switch using the BLS461 curve based on a 461 bit prime, hence upgrading our implementation to 128 bit security [15] against attacks looking for discrete logs on elliptic curves [16].
The future growth of quantum-computing technologies may be able to overcome the Discrete Logarithmic assumptions by qualitatively different computational means and SD-BLS may then be vulnerable to quantum-computing attacks. However this is speculative reasoning on what we can expect from the future.
The SD-BLS implementation we present in this paper doesn’t integrate any protection against replay-attacks, nor implements timestamps and expiration. We believe security features protecting from replay-attacks can be easily added and adapted from the context of deployment. I absence of session management providing such protections, any real-world use of SD-BLS should contemplate adding a signed timestamp to presentations and expiration checks on verifications.
5.2 Future development
In this section we describe possibilities for expanding the algorithm to cover further applications, which appear promising while requiring further investigation.
Compatibility with EUDI-ARF
EUDI-ARF dictates that the holder’s secret key generation and signatures must occur inside a trusted platform module (TPM). For mobile devices, this limits the secret keys and signatures to those offered, via proprietary APIs by the mobile OS, namely RSA (multiple flavours) and ECDSA on the secp256r1 curve. Currently the TPMs APIs supported by Android and iOS do not support BLS 12381 key generation or signature.
Client-side signatures in EUDI-ARF are mostly used in the authentication process, specifically in the proof of possession required by the OpenID4VCI[17] issuance flow, but not in the verification.
Therefore, we can investigate the possibility to use SD-BLS to implement a partially retro-compatible superset of EUDI-ARF, by maintaining the current issuance and verification protocols and using an extended SD-JWT format. Also we can explore useful integrations with the European Blockchain Services Infrastructure (EBSI [18]).
Signroom and DIDroom
In SD-BLS both credential issuers and revocation issuers are in charge of various interactive administrative operations, while the dealing of revocation shares can be easily automated.
We plan to integrate SD-BLS in the free and open source software “Signroom”, an application we developed in the context of the NGI ASSURE grant, and “DIDroom” the dashboard connected to our did:dyne W3C DID domain.
Digital Product Passport
Efforts in standardization of Digital Product Passport (DPP) are ongoing in both the EU (Cirpass, BatteryPass, Trace4EU) and US (DSCSA). An obstacle to adoption of DPP technologies is the reluctancy of manufacturers to share information about their supply-chain, knowing that the information would become publicly available and immutable due to blockchain storage. While requiring further analysis and investigation, a further development of the SD-BLS scheme could allow creating DPPs built on the selective disclosure principles, which may facilitate the adoption of the technology in the industry by preserving the privacy of natural persons present in the DPP as REA agents[19], while authenticating their contribution.
DAO Technologies
The SD-BLS math is fully compatible with ETH2.0 and can be computed inside an Ethereum VM. A verifier implemented in solidity can be a building block for more advanced Distributed Autonomous Organizations (DAO [20]) that want to authenticate peers using the selective disclosure of verifiable credentials instead of a key based proof of possession.
6 Acknowledgements
This work has been funded by the EU in the framework of the NGI TRUSTCHAIN project, grant No 101093274. We thank Puria Nafisi Azizi and Matteo Cristino for their infatiguable work on Zenroom, Luca Di Domenico for his help on the PVSS implementation, Giuseppe De Marco for sharing findings from his pioneering journey, Simone Onofri for his insights on threat models and Gabriele Bellini for his helpful security review. We extend our thanks the anonymous reviewers for their valuable advice and the De Cifris association for facilitating connections over a large network of unique professionals.